OpenStudy (anonymous):

Given the equation y − 4 = (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −x + y = 10 3x − 4y = −40 y = x + 12 y = x + 10

OpenStudy (anonymous):

Correction: y-4= 3/4(x+8

OpenStudy (anonymous):

)

OpenStudy (anonymous):

part c

OpenStudy (anonymous):

y=x+12, 4 moves to right